### v1

# LLMs Meet Long Video: Advancing Long Video Comprehension with An Interactive Visual Adapter in LLMs

Abstract
Long video understanding is a significant and ongoing challenge in the intersection of multimedia and artificial intelligence. Employing large language models (LLMs) for comprehending video becomes an emerging and promising method. However, this approach incurs high computational costs due to the extensive array of video tokens, experiences reduced visual clarity as a consequence of token aggregation, and confronts challenges arising from irrelevant visual tokens while answering video-related questions. To alleviate these issues, we present an Interactive Visual Adapter (IVA) within LLMs, designed to enhance interaction with fine-grained visual elements. Specifically, we first transform long videos into temporal video tokens via leveraging a visual encoder alongside a pretrained causal transformer, then feed them into LLMs with the video instructions. Subsequently, we integrated IVA, which contains a lightweight temporal frame selector and a spatial feature interactor, within the internal blocks of LLMs to capture instruction-aware and fine-grained visual signals. Consequently, the proposed video-LLM facilitates a comprehensive understanding of long video content through appropriate long video modeling and precise visual interactions. We conducted extensive experiments on nine video understanding benchmarks and experimental results show that our interactive visual adapter significantly improves the performance of video LLMs on long video QA tasks. Ablation studies further verify the effectiveness of IVA in long and short video understandings.
1 Introduction
The exponential advancement of the Internet and multimedia technologies has resulted in a significant surge in video content production by individuals and enterprises across various domains. The ability to interpret and extract meaningful content from videos is increasingly vital for meeting human demands and promoting the speed of information dissemination Tang et al. (2023). Therefore, Video Question Answering Yu et al. (2019); Li et al. (2023b); Castro et al. (2022) (Video QA), which allows users to ask about the content of videos through natural language and receive answers derived from their visual and auditory content, attracts tremendous research interest. Recently, large language models (LLMs) OpenAI (2023); Chiang et al. (2023) have demonstrated exceptional efficacy in the domains of human-machine interaction and the handling of extensive contextual information. Capitalizing on these advancements, there is a burgeoning inclination towards integrating LLMs into the realm of video information processing. This approach primarily aims to enhance the interface between users and video content through intelligent question-and-answer sessions.
The core of this innovation is a strategy that bridges the gap between the visual information in videos and the textual comprehension capabilities of LLMs. This is accomplished through a meticulously designed process that translates video data into a format comprehensible by LLMs, thereby facilitating an advanced question-answering system tailored for video content. The process involves mapping video encoding into the language space of LLMs via a learnable visual mapping network Wu et al. (2023); Li et al. (2023d); Dai et al. (2023). Essentially, the video is converted into “video tokens”, which are then fed into the LLM along with textual tokens of natural language questions. Leveraging the vast knowledge storage and natural language processing prowess of LLMs, this approach effectively handles video QA tasks. For instance, Maaz et al. (2023) performs spatial and temporal pooling for video tokens and feeds them into Vicuna Chiang et al. (2023) to achieve the interaction between users and video content. Zhang et al. (2023b) utilizes Q-former Li et al. (2023d) to extract question-relevant video tokens, which are then fed into LLama Gao et al. (2023) to generate the answer.
These LLMs-powered video understanding models Tang et al. (2023); Song et al. (2023) mainly focus on short video modeling and have achieved a successful performance on short video captioning Zhang et al. (2023c), question-answering Jin et al. (2023), and summarization Tang et al. (2023). However, the core challenges of video processing  Xu et al. (2023) stem from the need to efficiently model long video sequences and precisely respond to questions relevant to the video. Generally, using LLMs to handle long-form video often encounters the following hurdles: 1) high computational costs from a multitude of video tokens; 2) reduced visual clarity as a consequence of token aggregation such as employing average or maximum representation pooling for visual frames; 3) irrelevant visual tokens leading to incorrect answers, notably when question-relevant information is embedded within long temporal cues. Hence, previous models struggle to handle long-form videos owing to the constrained input capacity for video tokens and the challenge of distilling question-relevant, fine-grained visual features during generation.
To alleviate these issues, we present a long video comprehension method for LLMs, named Interactive Visual Adapter (IVA) to achieve in-depth interactions between LLMs and video content. Specifically, we first use the pretrained visual encoder to obtain global and fine-grained frame representations. We construct the temporal video tokens by integrating the global features of frames with temporal video embeddings, which are obtained through a pretrained causal transformer. The whole set of temporal video tokens is fed into the LLM to attain a whole understanding of the video content. Additionally, we designed a parameters-sharing Interactive Visual Adapter (IVA) that contains an instruction-aware temporal frames selector and a spatial feature interactor. The selector is used to obtain question-relevant frames based on contextual query embeddings and global encodings of videos. The selected frames are then fed into the spatial interactor to engage with the contextual query embeddings, in which fine-grained representations of frames are used. By doing so, LLMs could achieve in-depth interaction with video content by applying IVA between different layers.
To verify the effectiveness of our method, we conducted extensive experiments on four long video QA and five short video understanding benchmarks. Experimental results indicate that IVA is capable of achieving effective interactions between LLMs and long or short videos. Our contributions are summarized as follows:
  • We analyze the challenges of modeling long videos for LLMs and propose an interactive visual adapter for LLMs to handle long videos. It realizes the in-depth interaction between LLMs and long videos based on efficient video tokens and the IVA mechanism.
  • The proposed IVA is capable of selecting relevant frames and interacting with their fine-grained spatial features through the internal selector and interactor, respectively. The IVA architecture is lightweight and designed to be shareable between layers of LLMs.
  • Experimental results show that LLMs with IVA could achieve powerful performances in understanding long videos. Ablation studies underscore the critical role and effectiveness of IVA, confirming its significant contribution to enhanced performance.
2 Related Work
Traditional Video Understanding Models The rapid development of deep learning methods possesses superior task-solving capabilities for video understanding. DeepVideo Karpathy et al. (2014) was the earliest method introducing a Convolutional Neural Network (CNN), for video understanding. Two-stream networks Feichtenhofer et al. (2016), then integrating Convolutional Neural Networks (CNNs) Feichtenhofer et al. (2016) and Improved Dense Trajectories (IDT) Li et al. (2021), enhanced motion analysis in video understanding. For long-form content, Long Short-Term Memory (LSTM) Yue-Hei Ng et al. (2015) networks were adopted, offering a robust solution for sequential data analysis over extended durations. Additionally, Temporal Segment Network (TSN) Wang et al. (2016) advanced long-form video understanding by segmenting videos for individual analysis before aggregating insights, enabling more nuanced interpretation. Meanwhile, 3D networks started another branch by introducing 3D CNN to video understanding (C3D) Tran et al. (2015). The introduction of Vision Transformers (ViT) Dosovitskiy et al. (2021); Arnab et al. (2021); Fan et al. (2021) promotes a series of prominent models Among the pioneering efforts in this self-supervised video training domain, VideoBERT Sun et al. (2019) leverages the bidirectional language model BERT Kenton and Toutanova (2019) for self-supervised learning from video-text data. This model, and others following the "pre-training and fine-tuning" paradigm, such as ActBERT Zhu and Yang (2020), SpatiotemporalMAE Feichtenhofer et al. (2022), OmniMAE Girdhar et al. (2023), showcase the diverse strategies developed to enhance video understanding. Notably, these models have set a foundation for advanced video-language models like maskViT Gupta et al. (2022), CLIP-ViP Xue et al. (2022), LF-VILA Sun et al. (2022), further pushing the boundaries of what’s achievable in action classification, video captioning, and beyond. The evolution from VideoBERT to more recent innovations like HiTeA Ye et al. (2023), and CHAMPAGNE Han et al. (2023) underscores the rapid advancement in this field.
Figure 1: The overview of our framework employing LLMs to handle long video. While producing video tokens, we combine the global features and aggregated fine-grained features to represent a frame, allocating two tokens for each frame. The casual transformer is used to capture temporal relationships across frames and its output will be spliced with spatial feature sequence. The IVA will be inserted between blocks of LLMs to incorporate fine-grained visuals based on an understanding of the long video tokens, text instructions, and query tokens.
LLMs for Video Understanding The recent advancement in large language models (LLMs), pre-trained on expansive datasets, has ushered in groundbreaking capabilities in in-context learning Zhang et al. (2023a) and long-form context modeling Lyu et al. (2023). This innovation has paved the way for integrating LLMs with computer vision technologies, exemplified by initiatives like Visual-ChatGPT Wu et al. (2023). These models transcend traditional boundaries by calling vision model APIs Qin et al. (2023), thereby addressing complex problems within the computer vision domain. Integrating language models with video understanding technologies Maaz et al. (2023); Zhang et al. (2023d); Li et al. (2023e); Xu et al. (2023); Song et al. (2023) enhances multimodal understanding, facilitating sophisticated processing and interpretation of the intricate interplay between visual and textual data. They leverage their extensive multimodal knowledge base and nuanced contextual understanding, mirroring a more human-like comprehension of visual content. Moreover, the exploration of LLMs in video understanding tasks Tang et al. (2023) represents a significant stride towards harnessing their potential in analyzing and reasoning about visual data.
Multimodal Instruction Tuning for LLMs Recent advancements have significantly enhanced the performance of instruction-tuned, text-only large language models (LLMs) Ouyang et al. (2022); Muennighoff et al. (2022); Li et al. (2023g) on various NLP tasks and human-machine interaction scenarios. SFT LLMs have demonstrated remarkable capabilities in these areas. Building on this foundation, researchers are now exploring the integration of multimodal instruction data to further refine pre-trained LLMs, aiming to elevate their multimodal human-machine interaction competencies. For instance, a study Liu et al. (2023) employed GPT-4 to generate multimodal instruction data, which was then used to fine-tune the language model LLaMA on a synthetic dataset designed for instruction following. Similarly, another research Zhu et al. (2023) effort constructed a well-aligned multimodal instruction-following dataset to fine-tune Vicuna, an instruction-tuned language model, which showed superior performance in open-domain multimodal dialogues. Furthermore, a lightweight adaptation method Zhang et al. (2023d); Li et al. (2023h) was introduced to efficiently convert LLaMA into an instruction-following model, showcasing the potential for streamlined model enhancement. In this paper, we also explore the lightweight adapter to help LLMs understand long videos.
3 Methodology
3.1 Overview
Our work primarily introduces an interactive visual adapter for LLMs to handle long videos and answer relevant questions. The overview of workflow is shown in Figure 1. Specifically, given a video $V$, we first extract frames to obtain the whole sequence frame representations $\mathbf{h}_{V}=(\mathbf{h}_{I_{1}},...,\mathbf{h}_{I_{k}},...,\mathbf{h}_{I_{N}})$ via the pretrained image encoder, where $\mathbf{h}_{I_{k}}=(h_{g}^{I_{k}},h_{1}^{I_{k}}...,h_{576}^{I_{k}})$ refers to the representations of $k$ th frame and $N$ is the total number of extracted frames. Then, we use a casual transformer to acquire temporal video embeddings from the aggregated spatial representation. The overall video tokens are formed by merging temporal video embeddings and global spatial features $[h_{g}^{I_{1}},h_{g}^{I_{k}},...,h_{g}^{I_{N}}]$, where each frame is represented by two tokens. To enhance the capability of LLMs in leveraging fine-grained visual details from videos, we have developed an Interactive Visual Adapter (IVA) that is integrated into the blocks of LLM. This integration allows LLMs to comprehend the entirety of long videos through efficient video tokens while simultaneously capturing fine-grained visual information facilitated by the IVA.
3.2 Producing Video Tokens
We elaborate on the detailed process employed to produce efficient tokens for long videos, characterized by the extraction of one frame per second. First, we use the self-weighted calculation on the fine-grained feature $\mathbf{h}_{f}^{k}=(h_{1},...,h_{576})$ of a frame ($k$ th) to obtain its overall representation, which will be fed into the following casual transformer. This calculation process for the $k$ th frame is given in the following Eq. 1:
$\begin{array}[]{cc}s_{f}^{k}=Softmax(\mathbf{W}(\mathbf{h}_{f}^{k})+\mathbf{b}),\\ \mathbf{h}_{t}^{k}=s_{f}\mathbf{h}_{f}^{k},\end{array}$ (1)
where $s_{f}\in\mathbf{R}^{1\times 576}$ is the weight distribution and $\mathbf{W}$ and $\mathbf{b}$ are learnable parameters. Hence, we denote the obtained sequence-level frame representation as $\mathbf{h}_{f}=(\mathbf{h}_{f}^{1},...,\mathbf{h}_{f}^{N})$.
Casual Transformer is employed to acquire the temporal video embeddings. Specifically, we use a four-layer transformer to facilitate interaction across frames, where a frame only attends to its previous ones. Take the first layer as an example, the specific operation of the casual transformer is presented in Eq. 2:
$\begin{array}[]{cc}\mathbf{h}_{s}=SelfAtten(LayerN(\mathbf{h}_{f}),W^{Mask})+\mathbf{h}_{f}\\ \mathbf{h}_{s}=LayerN(\mathbf{h}_{s}),\\ \mathbf{h}_{o}^{1}=MLP(\mathbf{h}_{s})\end{array}$ (2)
where $SelfAtten$ and $LayerN$ are the self-attention calculation and the feature normalization. The top output of the casual transformer will be projected into the language model by a linear layer, which is spliced with the global features $\mathbf{h}_{g}^{V}=(h_{g}^{I_{1}},h_{g}^{I_{k}},...,h_{g}^{I_{N}})$. These global features of frames will be transferred into language models via a learnable MLP. We denote the final spliced feature to $\mathbf{h}_{V}=(h_{V}^{1},h_{V}^{2},...,h_{V}^{2N})$.
3.3 Interactive Visual Adapter
After obtaining video tokens $\mathbf{h}_{V}$, and supposing that the textual embeddings of instruction are initiated to $\mathbf{h}_{T}$ via the frozen word embedding table of LLMs, we concatenate them into a single sequence and fed it into LLMs. Considering fine-grained visual details existing in long videos, we expect that LMMs are capable of capturing the specific fine-grained visual information based on the understanding of instructions and the whole video representations. Hence, we devise a lightweight interactive visual adapter (IVA) to enable LLMs to focus on instruction-relevant fine-grained visuals during content generation.
Concretely, as the bottom part shown in Figure 1, we first introduce learnable dynamic tokens $\mathbf{h}_{D}=(h_{1}^{D},...,h_{M}^{D})$ as the query signals and integrate it at the end of the input token sequence. It aims to capture previous instruction and video information via the self-attention mechanism of LLMs, functioning as query tokens to engage with the fine-grained spatial features of videos. Suppose that the output of $i$ th layer of LLMs is $\mathbf{h}^{i}$. The specific calculation process of IVA between the $i$ and $i+1$ th layers of LLMs is shown in Eq. 3 and 4 in order. Each layer of IVA consists of a selector and an interactor, which are capable of selecting relevant frames and capturing valuable fine-grained visual information. The operational process of the selector is described as follows:
$\begin{array}[]{cc}\mathbf{h}^{S}_{q}=W^{q}\mathbf{h}^{i}_{d}+b^{q},\\ \mathbf{h}^{S}_{k}=W^{k}\mathbf{h}_{g}^{V}+b^{k},\\ \mathbf{M}^{S}=\mathbf{h}^{S}_{q}(\mathbf{h}^{S}_{k})^{T},\\ \mathbf{h}^{S}=Softmax(\mathbf{M}^{S}/\tau)Trans([\mathbf{h}_{I_{1}}^{f},...,\mathbf{h}_{I_{N}}^{f}]),\end{array}$ (3)
where $\mathbf{h}^{i}_{d}$ refers to the hidden states $\mathbf{h}^{i}$ associated with the indices of dynamic tokens. $W^{q}$, $W^{k}$, $b^{q}$, and $b^{k}$ are learnable parameters. $\mathbf{M}^{S}$ signifies the distribution score on the frames, which represents the relevant attention distribution. $\tau$ is the hyperparameter, which is set to 0.5. “Trans” refers to the transportation of feature dimension. $[\mathbf{h}_{I_{1}}^{f},...,\mathbf{h}_{I_{N}}^{f}]$ represents the fine-grained features of the entire video. The output $\mathbf{h}^{S}\in\mathbf{R}^{b\times M\times 576\times d_{S}}$ will be fed into the following interactor as the key value, where $d_{S}$ represents the dimension of the selector.
For the interactor, the specific calculation progress could be given as Eq. 4.
$\begin{array}[]{cc}\mathbf{h}^{I}_{q}=W^{1}\mathbf{h}^{i}_{d}+b^{1},\\ \mathbf{h}^{I}_{k}=W^{2}\mathbf{h}^{S}+b^{2},\\ \mathbf{M}^{I}=\mathbf{h}^{I}_{q}(\mathbf{h}^{I}_{k})^{T},\\ \mathbf{h}^{S}_{c}=Softmax(\mathbf{M}^{I})(W^{3}\mathbf{h}^{S}+b^{3}),\\ \mathbf{h}^{S}=MLP(\mathbf{h}^{S}_{c})+\mathbf{h}^{S}_{c}\end{array}$ (4)
where $W^{1}$, $W^{2}$, $W^{3}$, $b^{1}$, $b^{2}$, and $b^{3}$ are learnable parameters. Overall, we use the same four-layer calculations of the above selector and interactor to facilitate that LLMs interact with fine-grained visual features.
3.4 Training
Stage 1: Pretraining. To endow video tokens with meaningful representation, we first train the casual transformer, linear layers, and other learnable parameters during video tokens production, on massive video-caption pairs from WebVid, a total of 703k video-caption pairs. We freeze the other parameters of the overall model during this process and do not introduce the IVA module.
Stage 2: Video Instruction Tuning. At this stage, the model is required to generate responses that align with various instructions. These instructions often involve complex visual comprehension and reasoning, rather than merely describing visual signals. Note that the conversation data $[Q_{1},A_{1},...,Q_{r},A_{r}]$ consists of multiple rounds.
$X^{r}_{T}=\begin{cases}Q_{1},&r=1\\ \text{Concat}(Q_{1},A_{1},...,Q_{r},A_{r}),&r>1\end{cases}$ (5)
where $r$ represents the round count. As shown in Eq. 5, when $r>$ 1, we concatenate the conversations from all previous rounds with the current instruction as the input for this round. The training objective remains the same as the previous stage. After this stage, the model will be capable of generating corresponding responses based on various instructions.
Method ActivityNet-QA Social-IQ 2.0 LifeQA WildQA
Accuracy score Accuracy score Accuracy score Accuracy score
Video-LLaMA Zhang et al. (2023b) 12.4 1.1 55.8 2.9 35.8 2.3 63.2 3.2
Video-Chat Li et al. (2023f) 26.5 2.2 - - - - - -
LLaMA-Adapter Zhang et al. (2023d) 34.2 2.7 - - - - - -
Video-ChatGPT Maaz et al. (2023) 35.2 2.7 57.5 3.3 33.9 2.6 58.0 3.3
Baseline (w/o IVA) 40.8 3.0 46.8 3.0 31.7 2.3 52.2 3.1
IVA (LQ=8, NI=8) 41.6 3.0 54.0 3.6 46.5 2.8 51.2 3.1
IVA (LQ=16, NI=8) 42.1 3.0 64.9 3.9 50.5 3.0 53.5 3.2
IVA (LQ=32, NI=8) 41.9 3.0 57.1 3.7 51.9 3.1 53.7 3.2
IVA (LQ=16, NI=4) 42.2 3.0 63.3 3.9 50.1 3.0 52.5 3.2
IVA (LQ=16, NI=16) 42.3 3.0 55.4 3.7 50.0 3.0 55.1 3.3
IVA (LQ=16, NI=8)-272K 46.8 3.1 68.0 4.0 48.1 2.9 50.9 3.1
Table 1: Comparison between different methods on 4 zero-shot long video QA datasets. LLM with IVA achieves best performance on long videos compared to baselines and strong video LLMs. “LQ” refers to the length of query tokens and “NI” represents the number of interactions between LLMs and IVA. “-272K” indicates that we introduce additional training data of long video datasets like LifeQA and Social-IQ based on the original short video data.
4 Experiments
4.1 Data sets
While training the casual transformer, we utilize 702 thousand video-text pairs derived from Valley Luo et al. (2023), sourced from WebVid Bain et al. (2021). For the Video Instruction Tuning stage, we first collect instructional datasets from three sources. This includes a 100K video-text instruction dataset from Video-ChatGPT Muhammad Maaz and Khan (2023), a 36K short video-text instruction dataset from Valley-Instruct-73k Luo et al. (2023), and a 34K multiple-choice QA dataset from NExT-QA Xiao et al. (2021).
Additionally, we assessed the generalization of IVA using long and short video benchmarks. Long video benchmarks typically are characterized by videos exceeding one minute in duration. We evaluated our model using four prominent long video evaluation benchmarks: ActivityNet-QA Yu et al. (2019), Social-IQ 2.0 Wilf et al. (2023), LifeQA Castro et al. (2020), WildQA Castro et al. (2022). For short video benchmarks, the duration of the videos is often measured in seconds. We evaluated our model against three notable short video evaluation benchmarks: MSVD-QA Xu et al. (2017), MSRVTT-QA Xu et al. (2017), and SEED-Bench Li et al. (2023c).
4.2 Baselines
We mainly compare our models with the following video LLMs that could be extended to handle long videos. Video-ChatGPT Muhammad Maaz and Khan (2023) encodes frames independently and generates frame-level embeddings. Subsequently, it employs average pooling to transform these embeddings into both temporal and spatial features. These temporal and spatial features are then concatenated to derive video-level features and are fed into the LLM. Video-LLaMA Zhang et al. (2023c) utilizes Vision-Language and Audio-Language to process video frames and audio signals separately. After fine-tuning on image instruction dataset and video instruction dataset, Video-LLaMA exhibited remarkable abilities in comprehending images and videos. Video-Chat Li et al. (2023f) leverages perception tools to convert videos into textual descriptions in real-time, and employs a foundation model named InternVideo to encode videos into embeddings. These textual descriptions and video embeddings are then processed by an LLM for multimodal understanding. LLaMA-Adapter Zhang et al. (2023d) is a lightweight adapter injected into the attention calculation of LLM, which could be used to handle videos, text, and image tasks.
Method MSVD-QA MSRVTT-QA SEEDAR SEEDAP SEEDPU
Accuracy score Accuracy score Accuracy Accuracy Accuracy
Valley - - - - 31.3 23.2 20.7
Video-LLaMA 51.6 2.5 29.6 1.8 - - -
LLaMA-Adapter 54.9 3.1 43.8 2.7 - - -
Video-Chat 56.3 2.8 45.0 2.5 34.9 36.4 27.3
Video-ChatGPT 64.9 3.3 49.3 2.8 27.6 21.3 21.1
Baseline (w/o IVA) 54.5 3.2 49.6 2.9 22.5 23.5 24.8
IVA (LQ=8, NI=8) 53.2 3.2 47.6 2.9 32.0 31.8 27.5
IVA (LQ=16, NI=8) 55.7 3.2 49.1 2.9 35.2 32.0 34.2
IVA (LQ=32, NI=8) 53.0 3.2 47.2 2.9 32.2 32.1 28.8
IVA (LQ=16, NI=4) 55.0 3.2 47.8 2.9 32.5 31.7 26.0
IVA (LQ=16, NI=16) 53.3 3.1 47.1 2.8 31.8 29.4 31.0
IVA (LQ=16, NI-8)-272K 58.6 3.2 50.2 2.9 32.2 30.0 31.6
Table 2: Comparison between different methods on 5 zero-shot short video QA benchmarks. Benchmark names are abbreviated due to space limits. MSVD-QAXu et al. (2017); MSRVTT-QAXu et al. (2017); SEEDAR: SEED-Bench Action RecognitionLi et al. (2023c); SEEDAP: SEED-Bench Action PredictionLi et al. (2023c); SEEDPU: SEED-Bench Procedure UnderstandingLi et al. (2023c).
4.3 Evaluation Metrics
For open-ended video QA tasks, we employ ChatGPT-Assistant to evaluate the performance following Video-ChatGPT Muhammad Maaz and Khan (2023). First, we input the question, the predicted answer, and the correct answer into ChatGPT. Second, we request ChatGPT to verify the accuracy of the predicted answer, expecting a binary response of ’yes’ for correct predictions or ’no’ for incorrect ones. Additionally, we require ChatGPT to rate the quality of the predicted answer on a scale from 0 to 5, where 5 indicates a perfect match. Finally, we determine the overall accuracy by counting the number of ’yes’ responses and calculate the overall score by averaging all quality scores. This evaluation employs the "gpt-3.5-turbo" version of ChatGPT.
4.4 Implementation Details
We employ the AdamW optimizer Kingma and Ba (2014) in conjunction with a cosine learning rate scheduler to train our model. We first utilize 2 A100 GPUs to train visual-language MLP with 2 million image-text pairs with a global batch size of 256 and a base learning rate of 2e-4. Subsequently, we train the causal transformers using 703K video-text pairs data on the same two GPUs, employing a global batch size of 24 and a base learning rate of 3e-4. Transitioning to the video instruction tuning stage, we scale up to 8 A100 GPUs with a global batch size of 64. Here, we leverage LoRA to efficiently fine-tune the language model LLaMA. In our implementation, we set the rank to 128 and alpha to 256, maintaining a learning rate of 1e-4 for both LoRA and IVA parameters. Given the pretraining visual-language MLP and causal transformers, we adopt a smaller learning rate of 2e-5.
4.5 Main Results
We present the performance of the models on four long video QA benchmarks and five short video QA benchmarks. In zero-shot long video QA benchmarks, our model achieved state-of-the-art (SOTA) results compared to the previous pure video LLMs, except WildQA. Especially on the LifeQA and Social-IQ 2.0 evaluation datasets, our model achieved significantly higher results, surpassing the previous SOTA accuracy by 18.0 and 7.4 percentage points, respectively. In zero-shot short video QA benchmarks, our model also demonstrated strong capabilities across some evaluation datasets, especially in procedure understanding. Overall, IVA significantly enhances the capability of LLMs to analyze and interpret long videos, maintaining high-performance levels without compromising the understanding and reasoning abilities of short videos.
4.6 Ablation Study
Effect of IVA. From the results in Tables 1 and 2, we found that introducing the IVA module improved the overall visual understanding of the long video datasets of Social IQ2, LifeQA, and ActivityNet-QA, as well as the short video datasets. Among them, our model achieved an improvement of over 20% on LifeQA compared to the baseline, notably suggesting the effectiveness of IVA.
Length of Query Tokens. Comparing the experimental results of IVA (LQ=8, NI=8) and IVA (LQ=16, NI=8) in Tables 1 and 2, we observed a significant decrease in evaluation results across various benchmarks when reducing the length of query tokens (16 $\rightarrow$ 8). Regarding the comparison between IVA (LQ=16, NI=8) and IVA (LQ=32, NI=8) in long video benchmarks, we noted a slight decrease in performance on the first two benchmarks when increasing the length. However, while there was a slight improvement in LifeQA, it did not conclude an overall performance enhancement. In contrast, in the short video benchmarks, there was a downward trend in results across all benchmarks. Overall, increasing the length of query tokens may not lead to performance improvement. Moreover, reducing the length of query tokens may result in the loss of crucial visual information, consequently leading to performance degradation.
Figure 2: Five cases illustrate the comparative performances of our IVA Model and Baseline. The bottom part shows the detailed description of a video. Red words represent the inaccurate statement and the green words indicate the accurate statement.
Impact of long video instruction data. Furthermore, we attempted to further enhance the model’s performance during the Video Instruction Tuning phase by introducing more long video data. Therefore, we included the training sets of Social IQ2 and LifeQA, the video instruction part of the MIMIC-IT datasetLi et al. (2023a), along with the open-ended question-answering training data from NExT-QA, to the existing 170K training data, forming a new 272K training dataset. From the results of IVA(lQ=16, NI=8)-272K in the Tables 1 and 2, we observed a significant improvement on the long video dataset Social IQ2 with the inclusion of more training data. However, there was little difference in the results on the remaining datasets, and in some cases, a certain degree of decline was observed. This may be attributed to the somewhat indiscriminate addition of new datasets, leading to a certain imbalance in the proportions of different data. Additionally, the training set of LifeQA only consisted of 1,383 instances, which is relatively small in proportion to the total data, thus not providing sufficient improvement.
Number of Interaction between IVA and LLMs. Similar to the analysis of Query Tokens Length, we also conducted experiments with both doubled and halved number of interaction layers. The detailed injection layers are shown in Appendix A. Upon analyzing the results of IVA(LQ=16, NI=8) and IVA(LQ=16, NI=4) in Tables 1 and 2, we observe that this reduction resulted in a significant decrease in its performance on most long video datasets, especially on the Action Prediction and Procedure Understanding of the SEED-Bench. Moreover, the experimental results also indicate that increasing the number of layers (8 $\rightarrow$ 16) in the IVA interaction likewise caused a slight degradation in the model’s performance. Given that there was no significant improvement observed when increasing the interaction times between IVA and LLMs, we set it to 8 as the standard for experimentation.
4.7 Case Study
We present four open-ended question-answering cases and one detailed description example in Figure 2. Upon examining the initial two examples, we observe that the model augmented with IVA exhibits enhanced proficiency in recognizing particular actions associated with specific frames. In response to specific queries, it could discern objects such as the ’basketball-shaped cake’, which solely appears towards the video’s conclusion, and the ’glass bowl,’ present solely in the video’s opening segment. Furthermore, the fourth question-answering example illustrates that IVA augments the model’s reasoning ability, enabling it to deduce the prevailing weather conditions based on the lighting conditions within the video. These indicate the effectiveness of IVA in incorporating fine-grained visuals of long videos. Meanwhile, the bottom detailed description example reveals that when confronted with lengthy video descriptions, IVA could refine the perceptual acuity of LLMs, resulting in more precise recognition of elements such as the environment and color.
5 Conclusion
In this study, our primary goal is to enhance the capacity of LLMs to process and interpret long video content effectively. We identified the principal obstacles in this area and introduced an Interactive Visual Adapter (IVA) designed to facilitate dynamic interaction between LLMs and extended video sequences. The IVA incorporates a selector module for identifying relevant temporal frames within long videos based on specific instructions and tokens, along with an interactor module that isolates detailed spatial visual features within long videos. The empirical results demonstrate that our IVA significantly improves LLMs’ ability to comprehend and reason about long video content.
Limitations
Our work, while contributing valuable insights into video understanding through LLMs, is subject to several limitations that warrant further investigation:
  • Optimization for Longer Videos: Our current methodology demonstrates proficient performance in processing videos ranging from a few seconds to two minutes. However, the challenge of comprehensively understanding longer videos remains. Specifically, the optimization of video token length and the integration method of the Interactive Visual Adapter (IVA) within LLMs require further refinement to enhance their effectiveness and efficiency in handling extended content.
  • Impact of Interaction Frequency and Query Token Length: The stability of the IVA can be influenced by the frequency of interactions and the length of query tokens. These factors often occur in the development of multimodal large models, where a delicate balance must be struck between achieving high performance and maintaining operational efficiency, particularly in the context of long video interaction and encoding.
  • Accuracy and Appropriateness of Generated Responses: Another limitation is the potential for LLMs to generate responses that may be inaccurate, contain harmful content, or be factually incorrect. This issue stems from the inherent unpredictability in the response generation process of LLMs, underscoring the need for mechanisms that can ensure the reliability and appropriateness of the output.
Appendix A Inserting IVA in Different Layers
NI Corresponding Decoder Layers
4 0, 8, 16, 24
8 0, 4, 8, 12, 16, 20, 24, 28
16 0, 2, 4, 6, 8, 10, …, 22, 24, 26, 28, 30
Table 3: Ablation Study on Injection Layers for IVA. NI: Number of Inserting Layers. The incorporated inserting layers were positioned before the respective decoder layers.
In this section, we detail the methodology behind our ablation studies focusing on the variation in the Number of Injection Layers. Our experiments were structured around three different setups, where the injection layers were configured to be 4, 8, and 16 in number. To ensure a uniform distribution, these Injection Layers were interspersed throughout the decoder layers of the language model evenly. We utilized the Vicuna-7B model as our experimental framework, which is equipped with 32 decoder layers. The specific layers of the decoder that received the Injection Layers are outlined in Table 3, providing a clear reference to how the integration was achieved in each experimental setup